A company has just purchased a lift truck for £50,000 that has a useful life of...

  1. A company has just purchased a lift truck for £50,000 that has a useful life of 8 years. The engineer estimates that the maintenance costs by the end of the first year will be £1500 and are expected to increase as the truck ages at a rate of £200 annually over the remaining life. Also, the truck will need an overhaul that cost £5000 by the end of the 3rdand 6thyears. The firm decided to set up a maintenance account that earns 15 % to pay for all future maintenance expenses.
    1. Draw the cash flow for the truck costs.
    2. What is the maximum leasing value the company could make to lease a substitute truck?
    3. The lift’s work value is evaluated by £20,000; what is the payback period of the lift?
    4. Based on part (c) above, if the MARR of the company is 18%, does the lift satisfy the company standards? Comment.
Answer #1

a. Maintenance cost for

year 1 :1500

Year 2 : 1500+200

Year 3 : 1500+200+200

and so on

Overhaul cost is mentioned in the question for 3rd year and 6th year.

Purchase cost is allocated in the year 0 as given in the question.

Total cash outflow :

Year Maintenance Overhaul Cost Purchase Cost 50000 4 1900 5000 2500 6000 2100 2300 2900 0 0 1500 1700 0 0 2700 0 Total Outflows 50000 1500 1700 6900 2100 2300 8500 2700 2900 78600 Year Interest Work Value 315 345 225 20000 20000 20000 20000 20000 20000 20000 20000 0 255 285 375 405 435 Total Inflows 20225 20255 20285 20315 20345 20375 20405 20435 162640

c.

Payback period: the length of time required for an investment to recover its initial outlay in terms of profits or savings.

Initial Outlay: 78,600

60,765 will be recovered by the end of 3rd year.

Amount left to be recovered=  78,600- 60765 = 17835

Inflow in the fourth year = 20315

Payback period= 3 + 17835/20315 years

= 3 + 0.877 years

= 3.087 years


Year Total Inflows Cumulative flows 20225 20255 20285 20315 20345 20375 20405 20435 20465 20225 40480 60765 81080 4 9

d.

Since the NPV for the project will be positive, the lift does satisfy the company standards.

4 Year Maintenance Overhaul Cost Purchase Cost 50000 0OO 0 0 1500 17001900210023002500 27002900 0 0 5000 0 0 6000 0 0 Total Outflows 50000 1500 1700 6900 2100 2300 8500 2700 2900 PV Outflows 50000 1271.2 1220.9 4199.6 1083.2 1005.4 3148.668 847.6 771.5 63547.94 4 Year Interest Work Value 0 225 255285315345 20000 20000 20000 20000 200002000020000 20000 375 405 435 Total inflows PV Outflows 0 20225 20255 20285 20315 20345 20375 20405 20435 17140 14547 12346 10478 8893 7547.543 6406 5436 82793.64